OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. 4x^2 = 121 2.75, −2.75 5.5, −5.5 11, −11 30.25, −30.25

Nnesha (nnesha):

same like last one take square root both side or u can subtract 121 both side then do difference of square method

Nnesha (nnesha):

square root of 121 = ??

OpenStudy (anonymous):

11

OpenStudy (anonymous):

C

OpenStudy (hockeychick23):

its 5.5, -5.5

Nnesha (nnesha):

u have to take square root both side

Nnesha (nnesha):

http://openstudy.com/code-of-conduct DIRECT answer are not allowed

Nnesha (nnesha):

 u have to take square root both side

OpenStudy (anonymous):

ok thanks

Nnesha (nnesha):

\[\sqrt{4x^2} = \sqrt{4} \times \sqrt{x^2} = ???\]

Nnesha (nnesha):

 when u divide both side 2